# A Comparative Study of Competency Question Elicitation Methods from Ontology Requirements

### v1

@@ Abstract
Competency Questions (CQs) are pivotal in knowledge engineering, guiding the design, validation, and testing of ontologies. A number of diverse formulation approaches have been proposed in the literature, ranging from completely manual to Large Language Model (LLM) driven ones. However, attempts to characterise the outputs of these approaches and their systematic comparison are scarce. This paper presents an empirical comparative evaluation of three distinct CQ formulation approaches: manual formulation by ontology engineers, instantiation of CQ patterns, and generation using state of the art LLMs. We generate CQs using each approach from a set of requirements for cultural heritage, and assess them across different dimensions: degree of acceptability, ambiguity, relevance, readability and complexity. Our contribution is twofold: (i) the first multi-annotator dataset of CQs generated from the same source using different methods; and (ii) a systematic comparison of the characteristics of the CQs resulting from each approach. Our study shows that different CQ generation approaches have different characteristics and that LLMs can be used as a way to initially elicit CQs, however these are sensitive to the model used to generate CQs and they generally require a further refinement step before they can be used to model requirements.
Keywords:
Knowledge engineering Competency questions LLMs
@@ 1 Introduction
Competency Questions (CQs) [16] are fundamental in the ontology engineering lifecycle, as they are a way to bridge the gap between domain experts and knowledge engineers. Their primary function is to support the elicitation of the requirements, expressed in the form of questions in natural language, that an ontology should satisfy. CQs scope the knowledge modelled in an ontology and guide ontology design activities by suggesting relevant concepts and relations to model [30, 34]. CQs are also instrumental in other phases of the ontology construction process: they support ontology validation and testing by providing verifiable queries the ontology must answer [19, 8], and ontology reuse by clarifying the intended scope and capabilities of an existing ontology fragment [14, 6].
@@
Despite their recognised importance  [34, 29, 27], CQs are often underused in practice, for several reasons. Many developers find them difficult or too time-consuming to formulate, and lament the lack of standardised tool support [25, 3]. One further challenges to the effective adoption of CQs in practice lies in the absence of clear, objective criteria to assess whether generated CQs appropriately reflect intended requirements. As a result, practitioners often struggle to produce questions that are consistently clear, relevant, and suited to the task [21].
To facilitate the formulation and uptake of CQs, several research efforts have explored their (semi)-automatic generation, leveraging on controlled natural language, e.g. CLaRO [20], patterns and archetypes [33, 36, 8], and more recently, Large Language Models (LLMs) [5, 9, 31, 1, 37]. However, it remains unclear which generation methods are most effective, under what conditions, and for which user profiles, whether domain experts or ontology engineers, within the broader ontology development process.
This paper takes a first step towards addressing this challenge by investigating existing approaches for the formulation of CQs; we perform a comparative analysis of three state of the art approaches: (i) manual formulation by experienced ontology engineers; (ii) semi-automated formulation via instantiation of CQ patterns in [33]; (iii) automated formulation using LLMs (GPT 4.1, Gemini 2.5 Pro). In order to ensure consistency, the requirements are extracted from the same user story developed for a cultural heritage application.
@@
We systematically analyse the CQs produced by these approaches, using both a qualitative multi-annotator evaluation and a quantitative assessment to characterise the produced CQs in terms of their structural and linguistic features (readability and ambiguity, as well as their complexity and their relevance to the user story with respect to the underlying requirements). Overall, the paper presents the following contribution:
  • The first multi-annotator dataset of CQs generated from identical source material expressing ontological requirements from the user stories and using distinct elicitation approaches (manual, template, LLM-based).
@@
  • A framework for the comparative analysis of the CQs, which comprises: i. the evaluation of CQ suitability is manually performed by ontology engineers, ii. CQ feature extraction and analysis, to characterise the CQs generated by different approaches, and finally iii. the semantic overlap via embeddings to assess to what extent different approaches generate similar CQs.
Our results show that CQ formulation approaches have distinct characteristics: human formulated CQs have the highest acceptability score while producing the most readable and least complex questions. LLM based approaches generate a lower level of agreement, require more knowledge to be understood and are more complex, and therefore require further refinements.
@@
The remainder of this paper is structured as follows: Section 2 discusses related work, while Section 3 details the methodology used for the comparative study (the research questions, experimental design, and evaluation framework). We present the results in Section 4 and discuss key findings and limitation in Section 5. Finally, Section 6 concludes the paper and outlines future work.
@@ 2 Related work
Competency Questions (CQs) [35] are a core component of many ontology development methodologies and frameworks [27, 29, 34]. They support a range of ontology engineering activities, including scoping the ontology, guiding the reuse of existing ontological resources, informing the naming of classes and relations, and validating the resulting ontological artefacts [8, 12, 19, 22].
@@
CQs are widely acknowledged as an effective means of eliciting domain knowledge and aligning ontological models with stakeholder requirements. However, their practical application remains inconsistent and is at times intentionally bypassed [3, 25]. This is largely due to the lack of systematic methodological support for their formulation. Existing guidance is often fragmented and fails to clearly distinguish among the various purposes, stages, and quality criteria associated with CQ development. Despite these limitations, CQs continue to be regarded as a valuable asset for supporting ontology design and validation [21, 2].
@@
In recent years, a number of approaches have emerged that support users in eliciting CQs and therefore mitigate the CQ formulation issue. Ren and colleagues [33] propose a method to define ontology authoring tests through the definition of patterns (archetypes) that generate answerable CQs. The definition of patterns has also been proposed in [36], where the patterns are instantiated with domain knowledge to assist ontology engineers in formulating machine-processable CQs that can be used for ontology testing. Bottom-up approaches have also been proposed that use patterns for the generation of Controlled Natural Language (CNL) templates for authoring ontologies [20, 4]. However, both patterns and templates still require significant effort by ontology developers and domain experts since they are meant as a guide for formulating domain specific CQs across different scenarios [4].
The use of generative AI has further expanded the landscape of CQ authoring. Recent approaches leverage LLMs and external knowledge sources to support interactive CQ generation, often in conversational settings involving domain experts and ontology engineers. Examples include OntoChat [37], RevOnt [9], and the LLM-driven methods by Rebboud and colleagues and Alharbi and colleagues who generate CQs that are relevant for an existing ontology [31, 1].
@@
Despite these advances, a persistent gap remains in the evaluation of automatically or semi-automatically generated CQs. Current methods vary significantly in their assessment strategies. Some rely on similarity metrics (e.g., [1, 31]), while others incorporate human judgment [9, 5]. However, the scope of human evaluation is often narrow. For example, [9] focuses primarily on readability, whereas [5] considers a broader range of criteria including syntactic correctness, answerability, domain relevance, and coverage. Often these measures have been proposed for the assessment of automatic question generation approaches, typically complementing other characteristics such as: fluency, clarity, conciseness, informativeness, and complexity [26, 15, 28].
These evaluation metrics depend on the availability of ground truth, i.e. benchmarks of CQs formulated by ontology engineers for the construction of one or more ontologies [7, 13]. However, recent studies [17, 1, 18] have shown that even these benchmarks can suffer from poor or ambiguous phrasing, typos, or more fundamental issues such as unanswerable CQs.
@@ 3 Comparative analysis of CQ elicitation approaches
The efficacy of the selected CQ elicitation approaches (manual, template-based, and LLM-based) is assessed by performing a comparative analysis of the CQs they generate. The CQs in this study are all derived from the same user story-based requirements, therefore eliminating the risk of inconsistencies in the requirements. For this study, we investigate the following research questions:
@@
    Manual. Two human annotators (HA-1, HA-2 with at least 5 years of experience in OE) independently read the user story. Based on their understanding of the requirements expressed, each annotator formulated a set of CQs they deemed necessary to capture the scope and functionalities described.
@@
    Pattern-based. This set was generated by using a set of predefined 19 CQ templates (archetypes) from [33]. These templates (e.g., “Which [CE1] [OPE] [CE2]?”, “Is the [CE1] [CE2]? ”, “Which are [CE]? ”) were manually instantiated based on the entities, relations, and information needs identified within the user stories. This set was created by another ontology engineer with at least 5 years of experience in requirement engineering.
@@
The CQs in AskCQ are organised in five distinct sets: HA-1 (44), HA-2 (54), Pattern (38), GPT 4.1 (26), and Gemini 2.5 Pro (42). All generated CQs are assigned unique identifiers, and anonymised with respect to their generation method to ensure the evaluators were not biased in the subsequent evaluation phases. To the best of our knowledge, AskCQ is the first publicly available multi-annotator CQ dataset derived from manual, template-based, and LLM-based approaches applied to the same ontological requirements. The dataset and the user story are both publicly released under a CC-BY licence.
@@ 3.1 CQ-level suitability evaluation and feature extraction
The RQs outlined in Section 3 are addressed using a mixed-methods evaluation approach. For RQ1, we conducted an expert evaluation with ontology engineers to assess the suitability of each CQ within the user story context. For RQ2, both manual and NLP-based methods were used to extract and compare relevant CQ-level features. For RQ3, a semantic analysis based on sentence embeddings was employed to investigate the semantic overlap among the CQ sets.
3.1.1 Expert evaluation
@@
Each anonymised CQ was independently assessed by three annotators (N=3) with expertise in OE. Annotators were first required to review the user story and persona descriptions before being assigned to an evaluation pool. To mitigate potential bias, the evaluators were distinct from the individuals who manually generated the CQs for the same set. For each CQ, annotators were instructed to either accept or reject it based on its suitability for guiding OE tasks for the given user story. No specific quality criteria were provided, thereby safeguarding the annotators’ independent judgment and minimising evaluative bias, analogously to the way the elicitation process was conducted. Annotators were also given the opportunity to provide brief comments, particularly to justify rejections or to highlight specific concerns regarding a CQ. Comments that indicated ambiguity were subsequently identified through manual analysis and used to trigger further discussions and inform the corresponding evaluation metric.
3.1.2 CQ feature extraction
For each CQ we extracted 4 features: ambiguity, readability, relevance to the user story, and complexity. We describe their extraction methodology and rationale below.
Ambiguity
@@ 3.1.2 CQ feature extraction
is measured as a binary label attributed to a CQ based on the comments contributed in the expert evaluation, as CQs that were marked as ambiguous can still be considered suitable for the OE task. The same annotators who decided on the suitability reviewed these comments. Those CQs that received at least one comment triggered a discussion amongst the annotators, since the presence of the comment was interpreted as being indicative of an issue. After the discussion, the annotators finalised their suitability score of the CQ.
Relevance
@@
measures the degree to which a CQ aligns with the requirements expressed in the user story. This assessment was performed by an LLM (Gemini 2.5 Pro), which was instructed to rate the relevance of each CQ on a 4-point Likert scale. The scale was defined to capture different aspects of relevance from an ontology engineering perspective, where higher scores indicate closer alignment with explicitly stated or functionally necessary requirements. Specifically, the scale points represented the following: (4) The CQ addresses a requirement explicitly stated in the user story; (3) The CQ pertains to a requirement that, while not explicit, is inferable from the user story using domain knowledge and is functionally necessary for fulfilling the story’s goals; (2) The CQ is not directly inferable from the user story but holds some contextual relevance to the persona or their objectives within the story; (1) The CQ introduces an additional requirement that is neither expressed in, nor inferable from, the user story and is not considered necessary. A subset of these LLM judgments (12 CQs, with their expected ratings) were manually validated for prompt engineering.
@@
Flesch-Kincaid Grade Level [23] (FKGL) $15.5911.8\times(\frac{|Syl|}{|W|})+0.39\times(\frac{|W|}{|S|})-$ 15.59 Estimates the years of formal education (U.S. grade, from kindergarten/nursery to college) a person needs to understand a piece of writing on the first reading.
Dale-Chall Readability Index [10] (DCR) $0.1579\times(\frac{|DC_{DW}|}{|W|}\times 100)+0.0496\times(\frac{|W|}{|S|})$ DCI is a readability formula that estimates the comprehension difficulty of a text based on word familiarity and sentence structure. Unlike the other readability metrics, it uses a specific list of 3,000 familiar words that are easily understood by 4th-grade students.
Readability
Each CQ was assessed to gauge its ease of understanding. As done in [9], a suite of established readability indices designed to capture different aspects of textual difficulty were initially computed for each CQ using the textstat Python library. For the subsequent analysis and reporting presented in this paper, we selected the Flesch-Kincaid Grade Level (FKGL) and the Dale-Chall Readability Score (DCR) as representative readability features.
This selection was informed by a correlation analysis on AskCQ, which indicated that these two indices largely account for the variance observed across the broader set of metrics initially considered [^1] The correlation analysis is provided in the supplementary material.. The focus of the selected metrics is the estimated US grade level required for comprehension (FKGL) and the reliance on common vocabulary (DCR), as explained in Table 1. Generally, lower scores on such indices indicate better readability, suggesting that less specialist knowledge or lower formal educational attainment is required for comprehension.
@@
In addition, we also compute the length of each CQ in terms of the number of characters ($c_{0}$ complexity) as a coarse-grained indicator of verbosity and potential elaboration. Overall, we expect these four dimensions – length, requirement, linguistic, and syntactic structure –- to provide complementary perspectives on CQ complexity. A CQ might be semantically complex (e.g., requiring navigation of intricate partonomy or causality relations) yet linguistically simple (e.g., “What caused this event?”), scoring high on requirement metrics but low on linguistic/syntactic ones. Conversely, a CQ might be ontologically straightforward but phrased using complex sentence structures, scoring high on syntactic metrics but low on semantic ones. While we do not provide a formal definition of CQ complexity and acknowledge the limitations of our approach (e.g., reliance on scoring mechanisms and primitive induction), it is designed to: (i) target aspects likely to correlate with CQ complexity, and (ii) examine how these metrics vary across the different sets in AskCQ.

### v2

@@ Abstract
Competency Questions (CQs) are pivotal in knowledge engineering, guiding the design, validation, and testing of ontologies. A number of diverse formulation approaches have been proposed in the literature, ranging from completely manual to those that use Large Language Model (LLM). However, attempts to characterise the outputs of these approaches and their systematic comparison are scarce. This paper presents an empirical comparative evaluation of three distinct CQ formulation approaches: manual formulation by ontology engineers, instantiation of CQ patterns, and generation using state of the art LLMs. CQs are generated using each approach given a set of requirements for cultural heritage, and they are assessed empirically across different dimensions: degree of acceptability, ambiguity, relevance, readability and complexity. Our contribution is twofold: (i) the first multi-annotator dataset of CQs generated from the same source using different methods; and (ii) a systematic comparison of the characteristics of the CQs resulting from each approach. Our study shows that different CQ generation approaches have different characteristics and that LLMs can be used as a way to initially elicit CQs, however these are sensitive to the model used to generate CQs and they generally require a further refinement step.
@@ 1 Introduction
Competency Questions (CQs) [16] are fundamental in the ontology engineering lifecycle, as they can bridge the gap between domain experts and knowledge engineers. Their primary function is to support the elicitation of the requirements, expressed in the form of questions in natural language, that an ontology should satisfy. CQs scope the knowledge modelled in an ontology and guide ontology design activities by suggesting relevant concepts and relations to model [30, 34]. CQs are also instrumental in other phases of the ontology construction process, as they support: ontology validation and testing by providing verifiable queries the ontology must answer [7, 17]; and ontology reuse by clarifying the intended scope and capabilities of an existing ontology fragment [6, 14].
@@
Despite their importance  [26, 29, 34], CQs are often underused in practice as developers find them difficult or too time-consuming to formulate, and lament the lack of standardised tool support [3, 24]. A further challenge to their effective adoption lies in the absence of clear, objective criteria to assess whether the generated CQs appropriately reflect the intended requirements. As a result, practitioners often struggle to produce questions that are consistently clear, relevant, and suited to the task [18]. To facilitate CQ formulation and uptake, several studies have explored their (semi)-automatic generation, leveraging controlled natural language, e.g., CLaRO[20], patterns and archetypes [7, 33, 36], and more recently, Large Language Models (LLMs) [2, 5, 8, 31, 37]. However, no comprehensive studies exist that systematically assess the efficacy of different generation methods under different conditions, user profiles and stakeholders (i.e. domain experts or ontology engineers) within the broader ontology development process.
@@
In this paper, we perform a comparative analysis of three state of the art approaches by systematically analysing the resulting CQs: (i) manual formulation by experienced ontology engineers; (ii) semi-automated formulation via instantiation of CQ patterns [33]; (iii) automated formulation using LLMs (GPT 4.1, Gemini 2.5 Pro). Each approach is used with a common set of requirements extracted from the same user story within the cultural heritage domain. The analysis includes both a qualitative multi-annotator evaluation and a quantitative assessment to characterise the produced CQs in terms of their structural and linguistic features (readability and ambiguity, together with complexity and their relevance to the user story with respect to the underlying requirements). The paper’s contribution is twofold:
@@
  • A multi-faceted comparative analysis of the CQs comprising: (i) the evaluation of CQ suitability, manually performed by ontology engineers; (ii) CQ feature extraction and analysis, to characterise the CQs generated by different approaches; and finally, (iii) the semantic overlap via embeddings to assess to what extent different approaches generate similar CQs.
  • The first multi-annotator dataset of CQs generated from identical source material expressing ontological requirements from the user stories and using distinct elicitation approaches (manual, pattern-based, LLM-based).
This is one of the first comparative analysis of manual, pattern-based, and LLM-based approaches from a controlled source (a single user story). We intentionally fixed the domain and source text to run a controlled comparison and isolate the effect of the CQ generation method. Our evaluation includes: (i) manual expert assessment; (ii) ambiguity annotation; and (iii) feature extraction across five datasets. These steps are sensitive to domain-specific terminology and narrative complexity, so we needed a single fixed reference input (the user story). Our results show that CQ formulation approaches have distinct characteristics: human formulated CQs generate the highest agreement while producing the most readable and least complex questions, whereas LLM based approaches generate lower agreement, require more knowledge to be understood and are more complex, generally requiring further refinements.
@@
Section 2 discusses related work, whilst the methodology used for the comparative study (the research questions, experimental design, and evaluation framework) are detailed in Section 3. We present the results in Section 4 and discuss key findings and limitation in Section 5, before concluding in Section 6.
@@ 2 Related work
Competency Questions (CQs) [35] are a core component of many ontology development methodologies and frameworks [26, 29, 34], supporting a range of ontology engineering activities, including scoping the ontology, guiding the reuse of existing ontological resources, informing the naming of classes and relations, and validating the resulting ontological artefacts [7, 12, 17, 21]. Although they are widely proposed as a means of eliciting domain knowledge and aligning ontological models with stakeholder requirements, their practical application remains inconsistent [3, 24], largely due to the lack of systematic methodological support for their formulation. Existing guidance is often fragmented and fails to clearly distinguish among the various purposes, stages, and quality criteria associated with CQ development. Despite these limitations, CQs continue to be regarded as a valuable asset for supporting ontology design and validation [1, 18].
@@
Various approaches have recently emerged to support users in eliciting CQs: Ren et.al. propose a method to define ontology authoring tests through the definition of patterns (archetypes) that generate answerable CQs [33]. Wiśniewski et. al. also proposed instantiating patterns with domain knowledge to assist ontology engineers in formulating machine-processable CQs that can be used for ontology testing [36]. Bottom-up approaches have been proposed that use patterns for the generation of Controlled Natural Language (CNL) templates for authoring ontologies [4, 20]. However, both patterns and templates still require significant effort by ontology developers since they need to be contextualised to the different scenarios [4]. The use of generative AI has further broadened the landscape of CQ authoring; recent approaches leverage LLMs and external knowledge sources to support automatic CQ generation, possibly in conversational settings involving domain experts and ontology engineers. Examples include OntoChat[37], LLM-driven methods for CQ generation, e.g. the work of Rebboud, Y., et al. [31] RevOnt[8], and RetrofitCQ [2].
@@
Despite this progress, a gap remains in the evaluation of (semi-)automatically generated CQs. Current methods vary significantly in their assessment strategies; some rely on similarity metrics [2, 31], while others incorporate human judgment [5, 8]. However, the scope of human evaluation is often narrow; for example, [8] focuses primarily on readability, whereas [5] addresses syntactic correctness, answerability, domain relevance, and coverage. Often these measures have been proposed to assess question generation approaches, typically complementing others e.g. fluency, clarity, conciseness, and complexity [15, 25, 27]. These evaluation metrics depend on the availability of ground truth, i.e. benchmarks of CQs formulated by ontology engineers for the construction of one or more ontologies [10, 13]. Yet recent studies [2, 19, 28] have shown that even these benchmarks can suffer from poor or ambiguous phrasing, typos, or more fundamental issues such as unanswerable CQs.
@@ 3 Comparative analysis of CQ elicitation approaches
We proposed a framework to compare selected CQ elicitation approaches (manual, pattern-based, and LLM-based).The CQs in this study are all derived from the same user story-based requirements, therefore eliminating the risk of inconsistencies in the requirements. We investigate the following research questions:
@@
    Manual. Two human annotators (HA-1, HA-2 with at least 5 years of experience in OE) independently read the user story. The annotators were explicitly instructed to formulate CQs that they considered necessary for modelling the scope and functionalities described in the user story, without any specific constraints on the number, phrasing, or style of the questions.
@@
    Pattern-based. This set was generated by another ontology engineer with at least five years of experience in requirements engineering, using a predefined set of CQ templates (archetypes) from [33]. These templates (e.g., “Which [CE1] [OPE] [CE2]?”, “Find [CE1] with [CE2]”, “Which are [CE]?”) were manually instantiated based on the entities, relations, and information needs identified in the user story. We generated CQs from the user story using all patterns proposed by Ren et al. [33]. Most patterns could be instantiated, but four main patterns produced no CQs because their structure does not match the entities or relations present in the user story.
@@
The CQs in AskCQ are organised in five distinct sets: HA-1 (44), HA-2 (54), Pattern (38), GPT 4.1(26), and Gemini 2.5 Pro(42). All generated CQs are assigned unique identifiers, and anonymised with respect to their generation method to ensure the evaluators were not biased in the subsequent evaluation phases. To the best of our knowledge, AskCQ is the first publicly available multi-annotator CQ dataset derived from manual, pattern-based, and LLM-based approaches applied to the same ontological requirements. The dataset and the user story are both publicly released under a CC-BY licence.
@@ 3.1 CQ-level suitability evaluation and feature extraction
The RQs outlined in Section 3 are addressed using a mixed-methods evaluation approach. For RQ1, we conducted an expert evaluation with ontology engineers to assess the suitability of each CQ within the user story context. For RQ2, both manual and NLP-based methods are used to extract and compare relevant CQ-level features. For RQ3, a semantic analysis based on sentence embeddings was employed to investigate the semantic overlap among the CQ sets.
@@
Each anonymised CQ was independently assessed by three annotators (N=3) with expertise in OE. Annotators were first required to review the user story and persona descriptions before being assigned to an evaluation pool. To mitigate potential bias, the evaluators were distinct from the individuals who manually generated the CQs for the same set. They were instructed to either accept or reject each CQ based on its suitability for guiding OE tasks for the given user story. No specific quality criteria were provided, thereby safeguarding the annotators’ independent judgment and minimising evaluative bias, analogously to the way the elicitation process was conducted. Annotators were also given the opportunity to provide brief comments, particularly to justify rejections or to highlight specific concerns regarding a CQ. Comments that indicated ambiguity were subsequently identified through manual analysis and used to trigger further discussions and inform the corresponding evaluation metric. Ultimately, we employed this manual evaluation of suitability to capture the expert judgment of ontology engineers regarding domain correctness, ensuring the CQs were grounded in practical modelling requirements.
As there is no consensus on a single definition of CQ quality, we selected a multi-faceted set of metrics – ambiguity, relevance to the user story, readability, and complexity – to characterise the CQs from complementary perspectives and enable their comparison. Each metric is described below:
Ambiguity:
@@
measured as a binary label attributed to a CQ based on the comments contributed in the expert evaluation, as CQs that were marked as ambiguous can still be considered suitable for the OE task. The same annotators who decided on the suitability reviewed these comments. Those CQs that received at least one comment triggered a discussion amongst the annotators, since the presence of the comment was interpreted as being indicative of an issue. After the discussion, the annotators finalised their suitability score of the CQ. Therefore, we operationalised “Ambiguity” as lack of consensus: a CQ was flagged as ambiguous if at least one annotator disagreed on suitability or explicitly commented on unclear meaning.
Relevance:
@@
measures the degree to which a CQ aligns with the requirements expressed in the user story. This assessment was performed by an LLM (Gemini 2.5 Pro), which was instructed to rate the relevance of each CQ on a 4-point Likert scale. The scale was defined to capture different aspects of relevance from an ontology engineering perspective, where higher scores indicate closer alignment with explicitly stated or functionally necessary requirements. Specifically, the scale points represented the following: The CQ addresses a requirement explicitly stated in the user story - 4; The CQ pertains to a requirement that, while not explicit, is inferable from the user story using domain knowledge and is functionally necessary for fulfilling the story’s goals - 3; The CQ is not directly inferable from the user story but holds some contextual relevance to the persona or their objectives within the story - 2; The CQ introduces an additional requirement that is neither expressed in, nor inferable from, the user story and is not considered necessary - 1. A subset of these LLM judgments (12 CQs, with their expected ratings) were manually validated for prompt engineering. We consider this relevance measure serves as a consistent proxy for expert judgment.
@@
Flesch-Kincaid Grade Level (FKGL) [22] $11.8\times(\frac{|Syl|}{|W|})+0.39\times(\frac{|W|}{|S|})-15.59$ Estimates the years of formal education (U.S. grade, from kindergarten/nursery to college) a person needs to understand a piece of writing on the first reading.
Dale-Chall Readability Index (DCR) [9] $0.1579\times(\frac{|DC_{DW}|}{|W|}\times 100)+0.0496\times(\frac{|W|}{|S|})$ Estimates the comprehension difficulty of text based on word familiarity and sentence structure (using a specific list of 3,000 familiar words that are easily understood by 4th-grade students).
Readability:
Each CQ was assessed to gauge its ease of understanding. As done in [8], a suite of established readability indices designed to capture different aspects of textual difficulty were initially computed for each CQ using the textstat Python library. For the subsequent analysis and reporting presented in this paper, we selected the Flesch-Kincaid Grade Level (FKGL) and the Dale-Chall Readability Score (DCR) as representative readability features that estimate the reader education level required to understand the text.
This selection was informed by a correlation analysis on AskCQ, which indicated that these two indices largely account for the variance observed across the broader set of metrics initially considered. [^1] The correlation analysis is provided in the supplementary material. The focus of the selected metrics is the estimated US grade level required for comprehension (FKGL) and the reliance on common vocabulary (DCR), as explained in Table 1. Generally, lower scores on such indices indicate better readability, suggesting that less specialist knowledge or lower formal educational attainment is required for comprehension.
@@
In addition, we also compute the length of each CQ in terms of the number of characters ($c_{0}$ complexity) as a coarse-grained indicator of verbosity. Overall, we expect these four dimensions – length, requirement, linguistic, and syntactic structure –- to provide complementary perspectives on CQ complexity. A CQ might be semantically complex (e.g., requiring navigation of intricate partonomy or causality relations) yet linguistically simple (e.g., “What caused this event?”), scoring high on requirement metrics but low on linguistic/syntactic ones. Conversely, a CQ might be ontologically straightforward but phrased using complex sentence structures, scoring high on syntactic metrics but low on semantic ones. Despite lacking a formal definition of CQ complexity and having some methodological limitations (e.g., reliance on scoring mechanisms and basic induction), our method aims to (i) capture aspects that are likely to correlate with CQ complexity and (ii) analyze how these metrics differ across the AskCQ sets.